Define drift velocityand derive an expression for it

Grade:12th pass

1 Answers

Rup
6 years ago
The avg value of velocity of the electron through which it moves towards anode when external electric field is applied on the conductor.For finding it Vavg = 0Vd= a×(tr) a=accelerationtr= relaxation time F= -eE E= applied fieldAlso F= maThen ma= -eEa= -(eE/m)trVd = a(tr)Vd=-(eE/m)trWhere -ve sign indicates that drift velocity is vector quantity and movement of electron is in opp direction of electric field